Two oxide of a metal contained 71.2% and 78.8% respectively. Show that this information illustration the law of multiple proportion


1
Expert's answer
2021-05-25T07:51:33-0400

full information is not given so , we will take an example to get the way to solve these types of problems

The percentage of metal in two metal oxides, A and B is 74.2% and 59%, respectively. Show that the law of multiple proportions is obeyed.

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Metal \ oxides & percentage \ of \ metal & percentage \ of \ oxygen\\ \hline
 A & 74.2 & 25.8 \\
 \hdashline
 B & 59 & 41
\end{array}"


In A, 74.2 g of metal combines with 25.8 g of oxygen


combines with 1 g of oxygen

="\frac{74.2}{25.8}=2.87g"


In B, 59 g of metal combines with 41 g of oxygen


And combines with 1 g of oxygen

="\frac{59}{41}=1.43g"



∴ The ratio by weight of metal combining with fixed mass of oxygen in A and B

=2.87:1.43=2:1


Hence, law of multiple proportions is obeyed